Brine containing 3 lbs./gal of salt enters a large tank at the rate of 2 gals / min and the mixture well stirred leaves at 1.5gal / min. If the tank contains initially 100 gal of water, with 4 lbs. of dissolved salt.


a ) Find the amount of salt in the tank at any time t in minutes .


b ) Find the amount of salt in the tank after 4 minutes.

1
Expert's answer
2022-03-07T17:27:04-0500

Solution;

Let y represent the amount of salt in the tank at time t, where t is given in minutes;

"\frac{dy}{dt}=R_{in}-R_{out}"

"R_{in}=\frac{3lbs}{gal}×\frac{2gals}{min}=6lbs/min"

Volume of mixture at time time t;

"V(t)=100+(2-1.5)t"

"V(t)=100+0.5t"

"R_{out}=1.5×\frac{y}{100+0.5t}=\frac{1.5y}{100+0.5t}"

Hence;

"\frac{dy}{dt}=6-\frac{1.5y}{100+0.5t}"

Rewrite;

"\frac{dy}{dt}+\frac{1.5y}{100+0.5t}=6"

The general solution of the equation is;

"y(t)=e^{-\int(x)dx}\int e^{\int p(x)dx}Q(x)"

"y(t)=e^{-\int \frac{1.5}{100+0.5t}dt}×6\int e^{\frac{1.5}{100+0.5t}dt}"

"y(t)=-6(100+0.5t)^{-1.5}\int(100+0.5t)^{1.5}dt"

"y(t)=-6(100+0.5t)^{-1.5}\frac{(100+0.5t)^{2..5}}{2.5}+C"

"y(t)=\frac{100+0.5t}{2.5}+C"

The initial conditions are;

"t=0,y=4"

"y(0)=4=\frac{100}{2.5}+C"

"C=4-40=-36"

Amount at time t is;

"y(t)=\frac{100+0.5t}{2.5}-36"

Amount after t=4;

"y(4)=\frac{100+0.5(4)}{2.5}-36"

"y(4)=4.8lbs"